OpenStudy (matlee):

9.86 yesterday , 9.73 today what is percatage decrease?

OpenStudy (matlee):

\[\left[\begin{matrix}9.86 & 9.73 \\ 100 & x\end{matrix}\right]\]

OpenStudy (matlee):

?

OpenStudy (photon336):

\[\frac{ 9.73 }{ 9.86 } = \frac{ x }{ 100 }\]

OpenStudy (matlee):

oh yeah i messed that up thank you

OpenStudy (matlee):

so 9.73*100 = 9.86x

OpenStudy (photon336):

no problem :)

OpenStudy (photon336):

yep :)

OpenStudy (matlee):

973=9.86x ?

OpenStudy (matlee):

Ok thank you

OpenStudy (photon336):

this will give you the % but it says percentage decrease so I think you would take your final answer and subtract it from 100

OpenStudy (matlee):

Yeah i got 98.6815415821501;98.681541582 so subratact from 100?

OpenStudy (matlee):

the second one lol

OpenStudy (photon336):

98.7-100

OpenStudy (photon336):

lol that's a long decimal

OpenStudy (matlee):

Ty 1.3%

OpenStudy (photon336):

yep

OpenStudy (matlee):

So this would be a negative 1.3 or postivie?

OpenStudy (photon336):

positive you can say it decreased by 1.3%

OpenStudy (matlee):

Ok thank you!

OpenStudy (photon336):

anytime